you are considering two cable networks for your television. tata sky charges rs. 120 per month regardless of the channels watched. hathway on the other hand charges rs. 1/ channel per month. your monthly channel demand is given by the equation cp 150-50 p, where p is the price per channel. (1) how many channels per month will you watch with each network? (ii) based on consumer surplus obtained, which provider should you choose

1
Expert's answer
2022-03-08T12:39:32-0500

Tata sky charges "rs. 120" per month regardless of the channels watched.


Hathway on the other hand charges "rs. 1" per channel per month.


Demand equation is given as:

"Cp=150-50p"

"P= rs. 1"

"Cp=150-50(1)"

"Cp=100"

Hence 100 channels watched per month.


Hathway charges will be "=100*1=rs. 100" per month


Tata sky charges "rs. 120" per month


Hence Hathway is the better option.